# RoboDuet: Whole-body Legged Loco-Manipulation with Cross-Embodiment Deployment

### v4

@@ Abstract
Fully leveraging the mobile manipulation capabilities of a quadruped robot equipped with a robotic arm is non-trivial, as it requires controlling all degrees of freedom (DoFs) of the quadruped robot to achieve effective whole-body coordination. In this letter, we propose a novel framework RoboDuet, which employs two collaborative policies to realize locomotion and manipulation simultaneously, achieving whole-body control through mutual interactions. Beyond enabling large-range 6D pose tracking for manipulation, we find that the two-policy framework supports cross-embodiment deployment, allowing for the use of different quadruped robots or various robotic arms. Our experiments demonstrate that RoboDuet achieves a 42.5% improvement in average success rate over the baseline in mobile manipulation tasks employing whole-body control. These policies also enable zero-shot deployment across different quadruped robots in the real world. To support further research, we provide open-source code and additional videos on our website: locomanip-duet.github.io.
@@
Legged robot, whole-body control, mobile manipulation, cross-embodiment, reinforcement learning.
@@ I Introduction
Mobile robots have increasingly been deployed to assist humans and demonstrated remarkable capabilities [1, 2]. Typically, these robots are equipped with wheeled mobile bases, making them less adaptable to diverse terrains and unable to adjust their base posture. This limitation has sparked interest in developing legged robots to undertake manipulation tasks, offering enhanced versatility and adaptability in diverse environments. By employing whole-body control in legged robots and robotic arms, it is possible to effectively overcome terrain constraints and significantly expand the manipulation workspace of the arms [3, 4, 5]. However, training a legged loco-manipulation robot to achieve whole-body control like humans, along with accurate pose tracking capabilities, presents a substantial challenge to researchers.
Figure 1: RoboDuet is a framework that affords mobile-manipulation and cross-embodiment deployment. Top row: cross-embodiment deployment. This feature enables control across six configurations generated by the permutation of two robotic arms and three quadruped robots, allowing for component replacement without the need to retrain the entire system. Middle row: mobile manipulation. From left to right, the robot walks to pick up a small ball on grass, grasps a doll from a high table in a café, grabs a bottle on lower stairs, and picks up a cup from an office desk. Bottom row: whole-body control. Given multiple target end-effector poses, the robot can adjust its entire body posture while moving to reach the desired targets and maintain stability.
@@
In awareness of these challenges, we introduce the RoboDuet: an integrated legged loco-manipulation framework tailored for large-range 6D pose tracking and whole-body contorl. As shown in Fig. 1, RoboDuet endows the policies with the capabilities of cross-embodiment deployment across different quadruped robots and robotic arms, mobile manipulation for diverse tasks, and robust whole-body control. These capabilities greatly enhance the system’s adaptability to different environments and reduce the need for retraining.
To simplify the whole-body control problem, we adopt a cooperative policy mechanism, achieved through the coordinated collaboration of a locomotion policy and an arm policy. The interaction between the locomotion policy and the arm’s actions exhibits a duet-like performance, where the locomotion policy utilizes the actions of the arm as guidance to adjust its posture, while the arm is complemented by the actions of the locomotion policy aiming to expand the robot’s workspace. The training process for RoboDuet is structured in two stages, as illustrated in Fig. 2. In stage 1, we develop the locomotion policy to endow the legged robot with essential mobility capabilities. Following stage 1, stage 2 involves training the arm policy that can coordinate with the locomotion policy. We argue that employing a two-stage training strategy enhances the stability of the training process, resulting in the acquisition of agile and large-range 6D pose tracking agents.
@@
  • We propose a framework that can simultaneously achieve robust locomotion and agile 6D end-effector pose tracking, thus capable of mobile manipulation.
  • Our framework effectively coordinates the legged robot and the robotic arm with two collaborative yet separated policies, introducing the ability of whole-body control and cross-embodiment deployment.
  • We conduct extensive simulation and real-world experiments to demonstrate the whole-body control, gait stability, and cross-embodiment ability of our framework.
@@ II Related Works
II-A Learning-based Locomotion
Learning-based algorithms, especially deep reinforcement learning (DRL), have significantly advanced the locomotion of quadruped robots [6, 7, 8, 9]. In contrast to their control-based counterparts, which require extensive engineering for accurate physical modeling of dynamics [10, 11], learning-based algorithms rely primarily on straightforward reward functions to develop robust locomotion policies. Physics simulators represented by IsaacGym [12] enable efficient data sampling and the acquisition of privileged observations in simulations. Techniques such as RMA[13] and domain randomization have effectively reduced the sim2real gap. Currently, learning-based locomotion rivals its traditional model-based control counterparts in adaptability to navigate difficult terrains, climb stairs [14, 15, 16], and perform parkour [17, 18].
II-B Whole-body Control for Legged Robot
Given the superior locomotive capabilities of quadruped robots compared to chassis, there is an emerging interest in developing whole-body control for integrating legged robots and robotic arms to finish mobile manipulation tasks. The current technological landscape features three primary strategies. The first involves control-based techniques like model predictive control (MPC), which require extensive engineering efforts and generally exhibit limited adaptability and robustness in complex environments [19, 20]. The second strategy adopts learning algorithms to generate high-level commands for legged robot or robotic arm, which are then translated into low-level joint control instructions based on built-in controllers or inverse kinematics (IK) [4, 5, 21, 22]. However, these approaches lack effective coordination between the legged platform and the arm or suffer from infeasible IK solutions, thus failing to maximize the potential for coordinating the pose of the legged robot to extend the operational range of the arm. The third approach leverages DRL to realize whole-body control. In earlier research, a unified policy is trained to control the entire system, achieving only end-effector position tracking [3]. A follow-up study extends this approach to 6D task-space pose tracking [23], but it is limited by the constraints of data collection. The most recent research leverages keypoint tracking to train a whole-body end-effector pose tracking policy, which requires an additional policy to provide locomotion capabilities [24]. Consequently, there is a clear need for innovative frameworks that are capable of harnessing the full locomotive advantages of quadruped robots while ensuring seamless coordination between the upper arm and the lower legged platform.
@@ III Methods
Figure 2: Overview of the framework. RoboDuet operates in two stages. In Stage 1, the loco policy is trained with the robotic arm fixed, enabling the quadruped robot to achieve robust locomotion. In Stage 2, the loco and arm policies are trained simultaneously in a cooperative manner. The loco policy from stage 1 is reused, but the original body orientation commands are replaced by guidance signals $a_{t}^{arm^{G}}$ generated by the arm policy, enabling coordinated whole-body control. During the transition between the two stages, reward adjustment is introduced to effectively leverage locomotion priors, facilitating seamless integration for whole-body mobile manipulation.
@@ III-A Cooperative Policy for Whole-body Control
RoboDuet consists of a loco policy for locomotion and an arm policy for manipulation. The two policies are harmonized as a whole-body controller. Specifically, the loco policy adjusts its actions accordingly by following instructions from the arm policy. For each policy, we implement reinforcement learning algorithms to maximize the discounted expected return $\mathbb{E}_{\pi_{\theta}}\left[\sum_{t=0}^{T-1}\gamma^{t}r_{t}\right]$ to find the optimal parameters $\theta$, where $r_{t}$ represents the reward at time step $t$, $\gamma$ is the discount factor, and $T$ is the maximum episode length. We utilize the Proximal Policy Optimization (PPO) [25]algorithm for training.
@@ III-A1 Loco policy
The goal of the loco policy $\pi_{loco}$ is to follow a target command $\mathbf{c}_{t}=(v^{\mathrm{cmd}}_{x},v^{\mathrm{cmd}}_{y},{\omega}^{\mathrm{cmd}}_{yaw},\phi^{\mathrm{cmd}}_{pitch},\phi^{\mathrm{cmd}}_{roll})$, where $v^{\mathrm{cmd}}_{x,y}=(v^{\mathrm{cmd}}_{x},v^{\mathrm{cmd}}_{y})$ are the desired linear velocity in base frame along x- and y- axes, ${\omega}^{\mathrm{cmd}}_{yaw}$ is the desired angular velocity in yaw axis, $\boldsymbol{\phi}^{\mathrm{cmd}}=(\phi^{\mathrm{cmd}}_{pitch},\phi^{\mathrm{cmd}}_{roll})$ denotes the desired pitch and roll angle of the base which plays an important role in achieving whole-body control and many downstream tasks. The observation of loco policy $o_{t}^{loco}$ contains leg states $s^{leg}_{t}\in\mathbb{R}^{26}$ (leg joint positions and velocities), base states $\boldsymbol{\phi}_{t}$ (roll and pitch angles), target commands $\mathbf{c}_{t}$, clock time $\mathbf{t}_{t}$, and last leg action $a^{leg}_{t-1}\in\mathbb{R}^{12}$. The leg action $a^{leg}_{t}$ represents a joint position offset that is added to the default joint position to specify the target position for twelve leg joint motors.
@@ III-A2 Arm policy
The goal of the arm policy $\pi_{arm}$ is to accurately track the 6D pose. The observations of arm policy $o_{t}^{arm}$ is composed of arm states $s^{arm}_{t}\in\mathbb{R}^{12}$ (arm joint positions and velocities), target end-effector pose $\chi_{t}\in\mathbb{R}^{6}$, base states $\boldsymbol{\phi}_{t}$, and last arm action $a^{arm}_{t-1}\in\mathbb{R}^{8}$. The actions of the arm policy consist of two parts: the first six actions $a^{arm^{J}}_{t}\in\mathbb{R}^{6}$ represent the target joint position offsets corresponding to six arm joint actuators. Then, it will be concatenated with the output of the loco policy to achieve synchronous control of the overall system. It should be mentioned that the position targets are tracked using a proportional-derivative controller. To expand the manipulation workspace with whole-body control, the rest part of the arm policy $a_{t}^{arm^{G}}=(a_{t}^{arm^{G}_{p}},a_{t}^{arm^{G}_{r}})\in\mathbb{R}^{2}$ is used to replace $\boldsymbol{\phi}^{\mathrm{\mathrm{cmd}}}$, providing additional degrees of freedom for end-effector tracking to cooperate with the loco policy. Due to the distinct focuses of the two policies, only essential information is shared to maintain a favorable cooperative relationship while mitigating disruptions caused by factors such as differences in task learning efficiency.
@@ III-B Two-stage Training
To ensure robust mobile manipulation, we employ a two-stage training strategy. Our method enables a seamless transition between these stages by maintaining consistent input and output dimensions for the policies throughout the entire process. Since the primary tasks differ across the stages—stage 1 focuses on locomotion, while stage 2 emphasizes manipulation—we introduce reward adjustment to integrate the latter without compromising the performance of the former.
@@ III-B1 Stage 1
During stage one, most components of the reward $r^{loco}$ utilized by the loco policy are identical to those described in [16], which can be referred to for more details. However, we remove the body height tracking component $r^{loco}_{\boldsymbol{h}_{z}^{\mathrm{cmd}}}$ to allow the body height to adjust adaptively according to the task. Due to the additional payload introduced by the arm, we incorporate an energy reward $r^{loco}_{energy}$ for the quadruped robot to facilitate smoother motion [26], as formulated below:
@@
where $\tau_{i}$ and $\dot{q}_{i}$ is the torque and velocity of the $i^{th}$ leg joint.
@@ III-B2 Stage 2
This stage aims to coordinate locomotion and manipulation to achieve whole-body large-range mobile manipulation. The arm policy is activated simultaneously with all the robotic arm joints. We adopt 6D target pose of end-effector as policy input. To eliminate the influence brought by body rotation [3], we similarly use a posture-independent spherical coordinate to define the target end-effector pose $\chi_{t}$. The target position of end-effector is represented by radius $l$, latitude $p$, and longitude $y$. To improve the accuracy of end-effector orientation tracking, we use euler angles $[roll,pitch,yaw]$ in Z-Y-X order for sampling, which can intuitively exclude many illegal postures, and convert them to included angle along each axis of the coordinate. The mathematical form of sampling can be expressed as follows:
@@ III-B3 Reward Adjustment
While the rewards in stage 1 drive the quadruped robot to achieve robust locomotion, the fixed foot placements are not conducive to the whole-body control required in stage 2, when the robot needs to adapt its body posture by altering foot placements. To address this, the raibert heuristic footswing tracking reward $r^{loco}_{\boldsymbol{s}^{\mathrm{cmd}}}$ is replaced by the hip joint constraint reward $r^{loco}_{hip}$. Additionally, the velocity tracking reward $r_{v^{\mathrm{cmd}}_{x,y}},r_{{\omega}^{\mathrm{cmd}}_{yaw}}$ is scaled to half its original value to accommodate different tasks. The reward adjustments are detailed in Table I. To achieve more coordinated whole-body control, we introduce $r^{arm}_{smooth}$ and $r^{arm}_{guide}$ to promote smoother arm motion and enhance the control of the body orientation.
@@ IV-B Algorithm Comparison
In order to examine the performance of our proposed method on whole-body control, 6D pose tracking, cross-embodiment, and mobile manipulation, we set up comparisons of the following algorithms:
  • Floating Base+IK: A floating base policy without whole-body control for the quadruped robot, combined with an IK solver [27] for arm.
@@
The first setting serves as a baseline to evaluate the solution stability of our method. The primary distinctions among the remaining four settings lie in whether they use a cooperative policy and a two-stage training approach, which are the two key components of our method. This comparison allows us to assess the effectiveness of each component.

### v5

# RoboDuet: Learning a Cooperative Policy for Whole-body Legged Loco-Manipulation

@@ Abstract
Fully leveraging the loco-manipulation capabilities of a quadruped robot equipped with a robotic arm is non-trivial, as it requires controlling all degrees of freedom (DoFs) of the quadruped robot to achieve effective whole-body coordination. In this letter, we propose a novel framework RoboDuet, which employs two collaborative policies to realize locomotion and manipulation simultaneously, achieving whole-body control through mutual interactions. Beyond enabling large-range 6D pose tracking for manipulation, we find that the two-policy framework supports zero-shot transfer across quadruped robots with similar morphology and physical dimensions in the real world. Our experiments demonstrate that RoboDuet achieves a 23% improvement in success rate over the baseline in challenging loco-manipulation tasks employing whole-body control. To support further research, we provide open-source code and additional videos on our website: locomanip-duet.github.io.
@@
Legged robot, whole-body control, loco-manipulation, reinforcement learning.
@@ I Introduction
Mobile robots have increasingly been deployed to assist humans and demonstrated remarkable capabilities [1, 2]. Typically, these robots are equipped with wheeled mobile bases, making them less adaptable to diverse terrains and unable to adjust their base postures. This limitation has sparked interest in developing legged robots to undertake manipulation tasks, offering enhanced versatility and adaptability in diverse environments. By employing whole-body control in legged robots and robotic arms, it is possible to effectively overcome terrain constraints and significantly expand the manipulation workspace of the arms [3, 4, 5]. However, training a legged loco-manipulation policy to achieve whole-body control, along with accurate 6D pose tracking capabilities, presents a substantial challenge to researchers.
Figure 1: RoboDuet is a framework that affords loco-manipulation and zero-shot transfer on morphologically and dimensionally similar quadruped robots. Top row: zero-shot transfer. This feature enables control across six configurations generated by the permutation of two robotic arms and three quadruped robots, allowing for the replacement of quadruped robots without the need to retrain the entire system. Middle row: loco-manipulation. From left to right, the robot walks to pick up a small ball on grass, grasps a doll from a high table in a café, grabs a bottle on lower stairs, and picks up a cup from an office desk. Bottom row: whole-body control. Given multiple target end-effector poses, the robot can adjust its entire body posture while moving to reach the desired targets and maintain stability.
@@
In awareness of these challenges, we introduce the RoboDuet: an integrated legged loco-manipulation framework tailored for large-range 6D pose tracking and whole-body contorl. As shown in Fig. 1, RoboDuet endows the policies with the capabilities of zero-shot transfer across quadruped robots with similar morphology and physical dimensions, loco-manipulation for diverse tasks, and robust whole-body control. These capabilities greatly enhance the system’s adaptability to different environments and reduce the need for retraining.
To simplify the whole-body control problem, we adopt a cooperative mechanism, achieved through the coordinated collaboration of a locomotion policy and an arm policy. The interaction between the locomotion policy and the arm’s actions exhibits a duet-like performance, where the locomotion policy utilizes the actions of the arm as guidance to adjust its posture, while the arm is complemented by the actions of the locomotion policy aiming to expand the robot’s workspace. The training process for RoboDuet is structured in two stages, as illustrated in Fig. 2. In stage 1, we develop the locomotion policy to endow the legged robot with essential mobility capabilities. Following stage 1, stage 2 involves training the arm policy that can coordinate with the locomotion policy. We argue that employing a two-stage training strategy enhances the stability of the training process, resulting in the acquisition of agile and large-range 6D pose tracking agents.
@@
  • We propose a novel RL-based whole-body control framework for quadrupedal loco-manipulation that enables 6D end-effector pose tracking during locomotion.
  • We introduce a cooperative mechanism that decouples multi-task learning, allowing the two policies to collaborate while maintaining focus and resulting in enhanced loco-manipulation capabilities compared to the unified policy.
  • Our approach achieves at least a 23% improvement in success rate for challenging mobile manipulation tasks over model-based controllers and demonstrates zero-shot transfer capabilities across quadruped robots with a 2.7 kg increase in weight.
@@ II Related Works
In recent years, the capabilities of legged robot locomotion have advanced rapidly, particularly in traversing complex terrains [6, 7], thereby facilitating their integration into diverse and challenging application domains. Furthermore, given the redundant degrees of freedom of quadruped robots compared to wheeled or tracked platforms, there are growing research efforts focused on whole-body control for legged robots with manipulators to achieve loco-manipulation capabilities. The current technological landscape features three primary strategies. The first approach leverages control-based techniques such as model predictive control (MPC), which explicitly consider the system’s kinematic and dynamic properties, making them well-suited for scenarios involving dense contact forces. However, these methods often entail substantial engineering effort and exhibit inherent limitations in adaptability and robustness when applied to unknown environments [8, 9, 10]. The second strategy adopts learning algorithms to generate high-level commands for legged robot or robotic arm, which are then translated into low-level joint control instructions based on built-in controllers or inverse kinematics (IK) [4, 5, 11, 12]. However, these approaches lack effective coordination between the legged platform and the arm or suffer from infeasible IK solutions, thus failing to maximize the potential for coordinating the pose of the legged robot to extend the operational workspace of the arm. The third approach leverages deep reinforcement learning (DRL) to realize whole-body control. In earlier research, a pipeline was developed to integrate a model-based manipulator with an RL-based locomotion policy, demonstrating robustness against external force disturbances induced by the manipulator [13]. However, the locomotion policy focused solely on resisting external disturbances, lacking the ability to achieve whole-body coordination. Later study proposed training a unified policy to control the entire system for achieving whole-body control, but this approach were limited to achieving only end-effector position tracking [3]. A follow-up study extends this approach to 6D task-space pose tracking [14], but it is limited by the constraints of data collection. The most recent research leverages keypoint tracking to train a whole-body end-effector pose tracking policy, which requires an additional policy to provide locomotion capabilities [15]. This indicates that the method lacks the ability to perform simultaneous mobility and manipulation, which is crucial for executing dynamic tasks such as door opening. Consequently, there is a clear need for innovative frameworks that can fully harness the locomotive advantages of quadruped robots while ensuring seamless coordination between the upper arm and the lower legged platform.
@@ III Methods
Figure 2: Overview of RoboDuet. In Stage 1, the loco policy is trained with fixed robotic arm, enabling the quadruped robot to achieve robust locomotion. In Stage 2, the loco and arm policies are trained simultaneously in a cooperative manner. The loco policy from stage 1 is reused, but the original body orientation commands are replaced by guidance signals $a_{t}^{arm^{G}}$ generated by the arm policy, enabling coordinated whole-body control. During the transition between the two stages, reward adjustment is introduced to effectively leverage locomotion priors, facilitating seamless integration for whole-body loco-manipulation.
@@ III-A Cooperative Policy for Whole-body Control
RoboDuet consists of a loco policy for locomotion and an arm policy for manipulation. The two policies are harmonized as a whole-body controller. Specifically, the loco policy adjusts its actions by following the instructions of the arm policy.
@@ III-A1 Loco policy
The goal of the loco policy $\pi_{loco}$ is to follow a target command $\mathbf{c}_{t}=(v^{\mathrm{cmd}}_{x},v^{\mathrm{cmd}}_{y},{\omega}^{\mathrm{cmd}}_{yaw},\phi^{\mathrm{cmd}}_{pitch},\phi^{\mathrm{cmd}}_{roll})$, where $v^{\mathrm{cmd}}_{x,y}=(v^{\mathrm{cmd}}_{x},v^{\mathrm{cmd}}_{y})$ is the desired linear velocity in base frame along x- and y- axes, ${\omega}^{\mathrm{cmd}}_{yaw}$ is the desired angular velocity in yaw axis, $\boldsymbol{\phi}^{\mathrm{cmd}}=(\phi^{\mathrm{cmd}}_{pitch},\phi^{\mathrm{cmd}}_{roll})$ denotes the desired pitch and roll angles of the base which plays an important role in achieving whole-body control and many downstream tasks. The observation of loco policy $o_{t}^{loco}$ contains leg states $s^{leg}_{t}\in\mathbb{R}^{24}$ (leg joint positions and velocities), base states $\boldsymbol{\phi}_{t}$ (roll and pitch angles), target commands $\mathbf{c}_{t}$, clock time $\mathbf{t}_{t}$, and last leg action $a^{leg}_{t-1}\in\mathbb{R}^{12}$. The leg action $a_{t}^{leg}$ represents joint position offsets that are added to the default joint positions to specify the target positions for the twelve leg joint motors.
@@ III-A2 Arm policy
The goal of the arm policy $\pi_{arm}$ is to accurately track the 6D end-effector pose. The observation of arm policy $o_{t}^{arm}$ is composed of arm states $s^{arm}_{t}\in\mathbb{R}^{12}$ (arm joint positions and velocities), target end-effector pose $\chi_{t}\in\mathbb{R}^{9}$, base states $\boldsymbol{\phi}_{t}$, and last arm action $a^{arm}_{t-1}\in\mathbb{R}^{8}$. The actions of the arm policy consist of two parts: the first six actions $a^{arm^{J}}_{t}\in\mathbb{R}^{6}$ represent the target joint position offsets corresponding to six arm joint actuators. Then, it will be concatenated with the output of the loco policy to achieve synchronous control of the overall system. Notably, the position targets are tracked using a proportional-derivative controller. To expand the manipulation workspace with whole-body control, the rest part of the arm policy $a_{t}^{arm^{G}}=(a_{t}^{arm^{G}_{p}},a_{t}^{arm^{G}_{r}})\in\mathbb{R}^{2}$ is used to replace $\boldsymbol{\phi}^{\mathrm{\mathrm{cmd}}}$, providing additional degrees of freedom for pose tracking to cooperate with the loco policy. Due to the distinct focuses of the two policies, only essential information is shared to maintain a favorable cooperative relationship while mitigating disruptions caused by factors such as differences in task learning efficiency.
@@ III-B Two-stage Training
To ensure robust loco-manipulation, we employ a two-stage training strategy. Our method enables a seamless transition between these stages by maintaining consistent input and output dimensions for the policies throughout the entire process. Since the primary tasks differ across the stages—stage 1 focuses on locomotion, while stage 2 emphasizes manipulation—we introduce reward adjustment to integrate the latter without compromising the performance of the former.
@@ III-B1 Stage 1
During stage one, most components of the reward $r^{loco}$ utilized by the loco policy are identical to those described in [16], which can be referred to for more details. However, we remove the body height tracking component $r^{loco}_{\boldsymbol{h}_{z}^{\mathrm{cmd}}}$ to allow the body height to adjust adaptively according to the task. The additional payload introduced by the arm increases the torque demands on the leg motors, which will result in motor overheating [14]. To mitigate this, we incorporate an energy reward $r^{loco}_{energy}$ for the quadruped robot to ensure the stability of instantaneous power [17], as formulated below:
@@
where $\tau_{i}$ and $\dot{q}_{i}$ are the torque and velocity of the $i^{th}$ leg joint.
@@ III-B2 Stage 2
This stage aims to coordinate locomotion and manipulation to achieve whole-body large-range loco-manipulation. The arm policy is activated simultaneously with all the robotic arm joints. We adopt 6D target pose of end-effector as policy input. To eliminate the influence brought by body rotation [3], we similarly use a posture-independent spherical coordinate to define the target end-effector pose $\chi_{t}$. The target position of end-effector is represented by radius $l$, latitude $p$, and longitude $y$. The target orientation is represented using a 6-D vector [18]. To improve the accuracy of end-effector orientation tracking, we use euler angles $[roll,pitch,yaw]$ in Z-Y-X order for sampling, which can intuitively exclude many illegal postures, and convert them to included angle along each axis of the coordinate. The mathematical form of sampling can be expressed as follows:
@@ III-B3 Reward Adjustment
While the rewards in stage 1 drive the quadruped robot to achieve robust locomotion, the fixed foot placements are not conducive to the whole-body control required in stage 2, when the robot needs to adapt its body posture by altering foot placements. To address this, we replace the raibert heuristic footswing tracking reward $r^{loco}_{\boldsymbol{s}^{\mathrm{cmd}}}$ with the hip joint constraint reward $r^{loco}_{hip}$. Additionally, the velocity tracking rewards, $r_{v^{\mathrm{cmd}}_{x,y}}$ and $r_{{\omega}^{\mathrm{cmd}}_{yaw}}$, are scaled to half of their original values to accommodate different tasks. The reward adjustments are detailed in Table I. To achieve more coordinated whole-body control, we introduce $r^{arm}_{smooth}$ and $r^{arm}_{guide}$ to promote smoother arm motion and enhance the control of the body orientation. After the reward adjustments, the arm policy and locomotion policy will be synchronized to start stage 2 training.
@@ IV-B Algorithm Comparison
In order to examine the performance of our proposed method on whole-body control, 6D pose tracking, cross-embodiment, and loco-manipulation, we set up comparisons of the following algorithms:
  • Floating Base (FB)+IK: A floating base policy without whole-body control for the quadruped robot, combined with an IK solver [19] for arm.
  • Adaptive Floating Base (AFB)+ID: An adaptive floating base policy coupled with an inverse dynamics based arm control strategy, corresponding to the built-in controllers of the Unitree Go1 and ARX5.
@@
The first setting serves as a baseline to evaluate the solution stability of our method. The second setting represents a more powerful real-world baseline, where the legged robot actively adjusts its posture in response to the interaction forces induced by the arm’s motion. The primary distinctions among the remaining four settings lie in whether they use a cooperative policy and a two-stage training approach, which are the two key components of our method. This comparison allows us to assess the effectiveness of each component.